OpenStudy (anonymous):

help please x+7y=0 and 2x-8y=22

OpenStudy (anonymous):

what are u looking for here?

OpenStudy (anonymous):

x and y value?

OpenStudy (anonymous):

solve each system by substitution

OpenStudy (jojokiw3):

x = -7y. 2(-7y) - 8y = 22. -14y - 8 y = 22. -22y = 22 y = -1. x + 7(-1) = 0. x - 7 = 0 x = 7.

OpenStudy (anonymous):

thank you so much